# A positive meshless finite difference scheme for scalar conservation laws with adaptive artificial viscosity driven by fault detection

### v1

@@ Abstract
We present a meshless finite difference method for multivariate scalar conservation laws that generates positive schemes satisfying a local maximum principle on irregular nodes and relies on artificial viscosity for shock capturing. Coupling two different numerical differentiation formulas and adaptive selection of the sets of influence allows to meet a local CFL condition without any a priori  time step restriction. Artificial viscosity term is chosen in an adaptive way by applying it only in the vicinity of the sharp features of the solution identified by an algorithm for fault detection on scattered data. Numerical tests demonstrate a robust performance of the method on irregular nodes and advantages of adaptive artificial viscosity. The accuracy of the obtained solutions is comparable to that for standard monotone methods available only on Cartesian grids.
@@ 1 Introduction
Meshless methods discretize partial differential equations on irregular nodes in space without creating a grid or mesh Suchde et al. [2022], which allows to distribute the degrees of freedom in the most flexible way and rapidly rearrange them if needed as time evolves, without meshing or re-meshing, which eliminates a major time consuming and often algorithmically difficult step in numerical simulations. Especially, meshless finite difference methods, based on polynomials or radial basis functions, see for example Oñate et al. [1996], Fornberg and Flyer [2015], Jacquemin et al. [2020], attract increasing attention of engineers and mathematicians as a general purpose numerical technique that combines the basic simplicity and high accuracy inherited from the classical finite difference method with full spatial flexibility.
@@
We test numerically three versions of our scheme (no artificial viscosity, constant or adaptive artificial viscosity) on irregular Halton nodes for three bivariate benchmark problems: inviscid Burgers’ equation with either smooth or discontinuous initial condition, and the rotating wave problem with a non-convex flux. In all cases solutions are non-oscillating as expected due to the positivity of the schemes. The shocks are reproduced by both methods that employ artificial viscosity, while the version with adaptive viscosity is less dissipative. The accuracy of our solutions is comparable to those obtained by classical monotone schemes on Cartesian grids, and mostly somewhat better, despite using irregular nodes.
We expect that our positive first order scheme may be complemented by high order distretizations in conjuction with limiters, ENO/WENO and further ideas known in the literature on conservation laws.
@@ 2 Positive meshless finite difference schemes for scalar conservation laws
Because of all these difficulties, sophisticated adaptive algorithms are needed in order to obtain satisfactory numerical approximations of challenging conservation law problems arising in applications. Therefore numerical methods capable to address these challenges must be in position to quickly refine, coarsen and rearrange the discretization of the spacial computational domain when time evolves. Even if mesh-based numerical techniques such as finite volume method or discontinuous Galerkin method possess a great deal of flexibility, they are often slowed down by the pitfalls of the meshing technology. Meshless methods, in particular meshless finite differences, that rely solely on nodes distributed irregularly in the space, without the need to connect them into meshes, provide a promising alternative.
@@
Let $X=\{{\bf x}_{1},\ldots,{\bf x}_{N}\}\subset\Omega$ be an arbitrary finite set of discretization nodes, and $\Delta t>0$ a step size in time. Note that the term ${\bf F}^{\prime}(u)\nabla_{\bf x}u$ in (3) for each fixed ${\bf x}$ is the directional derivative of $u$ at ${\bf x}$,
@@
$\Pi^{d}_{q}=\operatorname{span}\{{\bf x}^{\alpha}:\alpha\in{\mathbb{Z}}^{d}_{+},\;|\alpha|<q\},\quad{\bf x}^{\alpha}:=x_{1}^{\alpha_{1}}\cdots x_{2}^{\alpha_{d}},\quad|\alpha|:=\alpha_{1}+\cdots+\alpha_{d}.$
@@ 3 Minimal numerical differentiation formulas
$Df({\bf z}):=\sum_{\alpha\in{\mathbb{Z}}_{+}^{d},\,|\alpha|\leq k}c_{\alpha}({\bf z})\frac{\partial^{|\alpha|}f}{\partial x_{1}^{\alpha_{1}}\cdots\partial x_{d}^{\alpha_{d}}}({\bf z})\,$ (20)
@@
be a linear differential operator of order $k$ applied to a sufficiently regular function $f$ at a point ${\bf z}$. We consider the numerical approximation
$\hat{D}f({\bf z}):=\sum_{i=1}^{n_{{\bf z}}}w_{i}f({\bf x}_{i})\,$ (21)
of $Df({\bf z})$ defined as a linear combination of the values of $f$ on a finite set $X_{{\bf z}}$ of scattered nodes, where
$X_{{\bf z}}:=\{{\bf x}_{i}:\,i=1,\ldots,n_{{\bf z}}\}\subset{\mathbb{R}}^{d}.$
@@
If the weight vector ${\bf w}=[w_{i}]_{i=1}^{n_{{\bf z}}}$ is chosen such that (21) is exact for polynomials up to a certain order $q>k$, [^1] The existence of the solution obtained by imposition of the exactness conditions, which may not be achieved for some special geometry of the set $X_{{\bf z}}$, is guaranteed for example if $X_{{\bf z}}$ is unisolvent for the polynomial space of interest. then the following upper bound for the error holds [Davydov and Schaback, 2018, Theorem 7],
$|\hat{D}f({\bf z})-Df({\bf z})|\leq\sigma({\bf z},X_{{\bf z}},{\bf w})\,h_{{\bf z},X_{{\bf z}}}^{r+\gamma-k}|f|_{r,\gamma,\Omega_{{\bf z}}}\,,\quad f\in C^{r,\gamma}(\Omega_{{\bf z}}),\quad k\leq r\leq q-1,$ (22)
where $\Omega_{{\bf z}}$ denotes a domain containing the set $S_{z}:=\displaystyle{\bigcup_{i=1}^{n_{{\bf z}}}[{\bf z}\,,\,{\bf x}_{i}]}\,,$
$\sigma({\bf z},X_{{\bf z}},{\bf w}):=h_{{\bf z},X_{{\bf z}}}^{k-r-\gamma}\sum_{j=1}^{n_{{\bf z}}}|w_{j}|\|{\bf x}_{j}-{\bf z}\|_{2}^{r+\gamma}\,,\qquad h_{z,X_{{\bf z}}}:=\max_{1\leq i\leq n_{{\bf z}}}\|{\bf x}_{i}-{\bf z}\|_{2}\,,$
@@
$|f|_{r,\gamma,Q}\,:=\,\frac{1}{(\gamma+1)\cdots(\gamma+r)}\,\Big{(}\sum_{|\alpha|=r}{r\choose{\alpha}}\Big{|}\frac{\partial^{|\alpha|}f}{\partial x_{1}^{\alpha_{1}}\cdots\partial x_{d}^{\alpha_{d}}}\Big{|}^{2}_{0,\gamma,Q}\Big{)}^{1/2},\qquad|f|_{0,\gamma,Q}:=\sup_{{\bf x},{\bf y}\in Q\atop{\bf x}\neq{\bf y}}\frac{|f({\bf x})-f({\bf y})|}{\|{\bf x}-{\bf y}\|_{2}^{\gamma}},$
@@
and $C^{r,\gamma}(Q)$, $r\geq 0$, $\gamma\in(0,1]$, denotes the space consisting of all $r$ times continuously differentiable functions $f$ on $Q$ such that $|f|_{r,\gamma,Q}<\infty$. Note that the value of $\sigma({\bf z},X_{{\bf z}},{\bf w})$ does not depend on the size $h_{{\bf z},X_{{\bf z}}}$ of the neighborhood of ${\bf z}$, but only on the cardinality and the shape of $X_{{\bf z}}$, as well as on the choice of the weights. Therefore the convergence order of the approximation as $h_{{\bf z},X_{{\bf z}}}\to 0$ is determined by the power of $h_{{\bf z},X_{{\bf z}}}$ in (22) unless the sets $X_{{\bf z}}$ are in specific bad positions (for example, contained in zero sets of some polynomials of order $q$), that normally do not realize on irregular nodes.
As soon as the exactness for polynomials of order up to $q>k$ does not determine the weights $w_{i}$ uniquely, in particular when $n_{{\bf z}}>\dim\Pi^{d}_{q}={d+q-1\choose d}$, different strategies for a suitable selection of the weights can be considered. The obvious choice of minimizing $\sigma({\bf z},X_{{\bf z}},{\bf w})$ leads to the $\ell_{1}$-minimal formulas that can be computed by linear programming. A computationally more attractive alternative is to minimize the weighted $\ell_{2}$-seminorm
$|{\bf w}|_{2,r+\gamma}:=\Big{(}\sum_{i=1}^{n_{{\bf z}}}w_{i}^{2}\|{{\bf x}_{i}}-{{\bf z}}\|_{2}^{2(r+\gamma)}\Big{)}^{1/2},$
@@
$|\hat{D}f({\bf z})-Df({\bf z})|\leq\sqrt{n_{{\bf z}}}\,\rho({\bf z},X_{{\bf z}})\,h_{{\bf z},X_{{\bf z}}}^{r+\gamma-k}|f|_{r,\gamma,\Omega_{{\bf z}}}\,,\quad f\in C^{r,\gamma}(\Omega_{{\bf z}}),\quad k\leq r\leq q-1,$ (23)
@@
$\rho({\bf z},X_{{\bf z}}):=\inf\Big{\{}\sigma({\bf z},X_{{\bf z}},{\bf w}):\,{\bf w}\in{\mathbb{R}}^{n_{{\bf z}}},\;Dp({\bf z})=\sum_{i=1}^{n_{{\bf z}}}w_{i}p({\bf x}_{i})\;\text{ for all }p\in\Pi^{d}_{q}\Big{\}}.$
Despite the additional factor $\sqrt{n_{{\bf z}}}$ in (23) in comparison to the $\ell_{1}$-minimal formulas, both types of formulas behave similarly in numerical experiments, see further details in Davydov and Schaback [2018].
@@
In the case of Laplace operator $\Delta f=\frac{\partial^{2}}{\partial x_{1}^{2}}+\cdots+\frac{\partial^{2}}{\partial x_{d}^{2}}$ it is often attractive to enforce the positivity of the weight vector ${\bf w}$ in (21), by requiring that ${\bf x}_{1}={\bf z}$ and $w_{i}\geq 0$ for $i=2,\ldots,n_{{\bf z}}$. (Then $w_{1}<0$ because of the exactness for constant.) This for example helps to ensure that system matrices of the resulting meshless finite difference method for the Poisson equation are M-matrices, with various benefits for the numerical solution, see Seibold [2008] and references therein. In [Davydov and Schaback, 2018, Section 4.3] an error bound is obtained for positive formulas for elliptic differential operators of second order.
We generate inequality constrained $\ell_{2}$-minimal formulas in the form (5) and (12) for the operators $D_{{\boldsymbol{\eta}}_{i}}({\bf x}_{i})$ and $\Delta_{\bf x}({\bf x}_{i})$ as follows. The algorithm depends on the prescribed bounds $n_{\min},n_{\max}$ for the size of the set of influence and the artificial viscosity parameter $\mu_{i}\geq 0$. We start by choosing an initial set of influence $X_{i}=X_{i}^{init}$ consisting of $|X_{i}^{init}|=n_{\min}$ nearest neighbors of ${\bf x}_{i}$ in $X$, including ${\bf x}_{i}$ itself. (We use $n_{\min}=10$ in the numerical experiments in the bivariate case $d=2$.)
@@
and $X_{i}$ is initialized as $X_{i}^{init}$ if $\mu_{i}=0$ or $X_{i}^{visc}$ otherwise, and repeatedly expanded to the set of $\lceil 1.2\,|X_{i}|\rceil$ nearest neighbors of ${\bf x}_{i}$ as long as (25) is infeasible. Note that $v_{ii}$ is negative as soon as $\mu_{i}>0$ because of (17) and since $v_{ij}$, ${\bf x}_{j}\in X_{i}$, cannot all be zero thanks to (13). Again, we terminate the process if the new size is greater than $n_{\max}$, meaning the algorithm exits with failure. In this case we repeat the same procedure as above after removing the inequality constraints in (25) and resetting $X_{i}$ to $X_{i}^{init}$ if $\mu_{i}=0$ or $X_{i}^{visc}$ otherwise. With $n_{\max}=100$ in the bivariate case this happened very rarely and only for certain interior nodes situated very close to the inflow boundary nodes in Example 1 below, see Section 6.1. Note that the largest size of $X_{i}$ observed in the numerical experiments of Section 6 was 22 for both the viscosity term ($X_{i}^{visc}$) and the divergence term (final $X_{i}$).

### v2

@@ Abstract
We present a meshless finite difference method for multivariate scalar conservation laws that generates positive schemes satisfying a local maximum principle on irregular nodes and relies on artificial viscosity for shock capturing. Coupling two different numerical differentiation formulas and the adaptive selection of the sets of influence allows to meet a local CFL condition without any a priori  time step restriction. The artificial viscosity term is chosen in an adaptive way by applying it only in the vicinity of the sharp features of the solution identified by an algorithm for fault detection on scattered data. Numerical tests demonstrate a robust performance of the method on irregular nodes and advantages of adaptive artificial viscosity. The accuracy of the obtained solutions is comparable to that for standard monotone methods available only on Cartesian grids.
@@ 1 Introduction
Meshless methods discretize partial differential equations on irregular nodes in space without creating a grid or mesh Suchde et al. [2022]. This allows to distribute the degrees of freedom in the most flexible way and rapidly rearrange them if needed as time evolves, without meshing or re-meshing, and thus eliminating a major time consuming and often algorithmically difficult step in numerical simulations. Especially, meshless finite difference methods, based on polynomials or radial basis functions, see for example Oñate et al. [1996], Fornberg and Flyer [2015], Jacquemin et al. [2020], attract increasing attention of engineers and mathematicians as a general purpose numerical technique that combines the basic simplicity and high accuracy inherited from the classical finite difference method with full spatial flexibility.
@@
We test numerically three versions of our scheme (no artificial viscosity, constant or adaptive artificial viscosity) on irregular nodes for three bivariate benchmark problems: inviscid Burgers’ equation with either smooth or discontinuous initial condition, and the rotating wave problem with a non-convex flux. In all cases solutions are non-oscillating as expected due to the positivity of the schemes. The shocks are reproduced by both methods that employ artificial viscosity, while the version with adaptive viscosity is less dissipative. The accuracy of our solutions is comparable to those obtained by classical monotone schemes on Cartesian grids, and mostly somewhat better, despite using irregular nodes.
We expect that our positive first order scheme may be complemented by high order discretizations in conjuction with limiters, ENO/WENO and further ideas known in the literature on conservation laws.
@@ 2 Positive meshless finite difference schemes for scalar conservation laws
Because of all these difficulties, sophisticated adaptive algorithms are needed in order to obtain satisfactory numerical approximations of challenging conservation law problems arising in applications. Therefore numerical methods capable to address these challenges must be in position to quickly refine, coarsen and rearrange the discretization of the spatial computational domain when time evolves. Even if mesh-based numerical techniques such as finite volume method or discontinuous Galerkin method possess a great deal of flexibility, they are often slowed down by the pitfalls of the meshing technology. Meshless methods, in particular meshless finite differences, that rely solely on nodes distributed irregularly in the space, without the need to connect them into meshes, provide a promising alternative.
@@
Let $X=\{{\bf x}_{1},\ldots,{\bf x}_{N}\}\subset\overline{\Omega}$ be an arbitrary finite set of discretization nodes, and $\Delta t>0$ a step size in time. Note that the term ${\bf F}^{\prime}(u)\nabla_{\bf x}u$ in (3) for each fixed ${\bf x}$ is the directional derivative of $u$ at ${\bf x}$,
@@
$\Pi^{d}_{q}=\operatorname{span}\{{\bf x}^{\alpha}:\alpha\in{\mathbb{Z}}^{d}_{+},\;|\alpha|<q\},\quad{\bf x}^{\alpha}:=x_{1}^{\alpha_{1}}\cdots x_{d}^{\alpha_{d}},\quad|\alpha|:=\alpha_{1}+\cdots+\alpha_{d}.$
@@
Even if the choice of the time step $\Delta t$ and the viscosity coefficients $\mu_{i}$ are in general problem dependent, and fine tuning them may significantly improve the results in the applications, it is important to take into account their natural rescaling depending on the size of the characteristic velocity. It is easy to see that the solution of (3) becomes $u(\gamma t,{\bf x})$ if ${\bf F}^{\prime}(u)$ is replaced by $\gamma{\bf F}^{\prime}(u)$, and the solution of (11) becomes $u(\gamma t,{\bf x})$ if ${\bf F}^{\prime}(u)$ is replaced by $\gamma{\bf F}^{\prime}(u)$, while $\mu$ is replaced by $\gamma\mu$, for any $\gamma>0$. Therefore, in order to obtain the same behavior also in the discrete case, one needs to ensure that the discrete solution $U(t,{\bf x}_{i})$ becomes $U(\gamma t,{\bf x}_{i})$ under the same scaling of ${\bf F}^{\prime}(u)$ and $\mu_{i}$. Fortunately, this is easy to achieve for the scheme (14) by replacing the time step $\Delta t$ with $\gamma^{-1}\Delta t$, and $\mu_{i}$ with $\gamma\mu_{i}$, because the coefficients $w_{ij}$ will become $\gamma w_{ij}$, and hence $\gamma$ cancels out in both the divergence and the viscosity terms of (11). This justifies the usual approach of making $\mu$ directly and $\Delta t$ inversely proportional to the size of the characteristic velocity ${\bf F}^{\prime}(u)$.
@@ 3 Minimal numerical differentiation formulas
$Df({\bf x}):=\sum_{\alpha\in{\mathbb{Z}}_{+}^{d},\,|\alpha|\leq k}c_{\alpha}({\bf x})\frac{\partial^{|\alpha|}f}{\partial x_{1}^{\alpha_{1}}\cdots\partial x_{d}^{\alpha_{d}}}({\bf x})\,$ (20)
@@
be a linear differential operator of order $k$ applied to a sufficiently regular function $f$ at a point ${\bf x}$. For any ${\bf x}\in\overline{\Omega}$ and a set $X_{loc}\subset X$ of its neighbors in $X$, consider the numerical approximation
$\hat{D}f({\bf x}):=\sum_{{\bf x}_{j}\in X_{loc}}w_{j}f({\bf x}_{j})\,$ (21)
of $Df({\bf x})$ defined as a linear combination of the values of $f$ on $X_{loc}$. In this paper we use two types of operators: directional derivatives $D=D_{{\boldsymbol{\eta}}_{i}}$ and the Laplace operator $D=\Delta$, and perform numerical differentiation at nodes ${\bf x}={\bf x}_{i}$ on the influence sets $X_{loc}=X_{i}$ with weights denoted $w_{j}=w_{ij}$ for $D_{{\boldsymbol{\eta}}_{i}}$ and $w_{j}=v_{ij}$ for $\Delta$.
@@
If the weight vector ${\bf w}=[w_{j}]_{{\bf x}_{j}\in X_{loc}}$ is chosen such that (21) is exact for polynomials up to a certain order $q>k$, [^1] The existence of the solution obtained by imposition of the exactness conditions, which may not be achieved for some special geometry of the set $X_{loc}$, is guaranteed for example if $X_{loc}$ is unisolvent for the polynomial space of interest. then the following upper bound for the error holds [Davydov and Schaback, 2018, Theorem 7],
$|\hat{D}f({\bf x})-Df({\bf x})|\leq\sigma({\bf x},X_{loc},{\bf w})\,h_{{\bf x},X_{loc}}^{r+\gamma-k}|f|_{r,\gamma,\Omega_{loc}}\,,\quad f\in C^{r,\gamma}(\Omega_{loc}),\quad k\leq r\leq q-1,$ (22)
where $\Omega_{loc}$ denotes a domain containing the set ${\bigcup_{{\bf x}_{j}\in X_{loc}}[{\bf x}\,,\,{\bf x}_{j}]}\,,$
$\sigma({\bf x},X_{loc},{\bf w}):=h_{{\bf x},X_{loc}}^{k-r-\gamma}\sum_{{\bf x}_{j}\in X_{loc}}|w_{j}|\|{\bf x}_{j}-{\bf x}\|_{2}^{r+\gamma}\,,\qquad h_{{\bf x},X_{loc}}:=\max_{{\bf x}_{j}\in X_{loc}}\|{\bf x}_{j}-{\bf x}\|_{2}\,,$
@@
$|f|_{r,\gamma,\Omega_{loc}}\,:=\,\frac{1}{(\gamma+1)\cdots(\gamma+r)}\,\Big{(}\sum_{|\alpha|=r}{r\choose{\alpha}}\Big{|}\frac{\partial^{|\alpha|}f}{\partial x_{1}^{\alpha_{1}}\cdots\partial x_{d}^{\alpha_{d}}}\Big{|}^{2}_{0,\gamma,\Omega_{loc}}\Big{)}^{1/2},\quad|f|_{0,\gamma,\Omega_{loc}}:=\sup_{{\bf x},{\bf y}\in\Omega_{loc}\atop{\bf x}\neq{\bf y}}\frac{|f({\bf x})-f({\bf y})|}{\|{\bf x}-{\bf y}\|_{2}^{\gamma}},$
@@
and $C^{r,\gamma}(\Omega_{loc})$, $r\geq 0$, $\gamma\in(0,1]$, denotes the space consisting of all $r$ times continuously differentiable functions $f$ on $\Omega_{loc}$ such that $|f|_{r,\gamma,\Omega_{loc}}<\infty$. Note that the value of $\sigma({\bf x},X_{loc},{\bf w})$ does not depend on the size $h_{{\bf x},X_{loc}}$ of the neighborhood of ${\bf x}$, but only on the cardinality and the shape of $X_{loc}$, as well as on the choice of the weights. Therefore the convergence order of the approximation as $h_{{\bf x},X_{loc}}\to 0$ is determined by the power of $h_{{\bf x},X_{loc}}$ in (22) unless the sets $X_{loc}$ are in specific bad positions (for example, contained in zero sets of some polynomials of order $q$), that normally do not realize on irregular nodes.
[…]
$|{\bf w}|_{2,r+\gamma}:=\Big{(}\sum_{{\bf x}_{j}\in X_{loc}}w_{j}^{2}\|{{\bf x}_{j}}-{{\bf x}}\|_{2}^{2(r+\gamma)}\Big{)}^{1/2},$ (23)
@@
$|\hat{D}f({\bf x})-Df({\bf x})|\leq\sqrt{n_{loc}}\,\rho({\bf x},X_{loc})\,h_{{\bf x},X_{loc}}^{r+\gamma-k}|f|_{r,\gamma,\Omega_{loc}}\,,\quad f\in C^{r,\gamma}(\Omega_{loc}),\quad k\leq r\leq q-1,$ (24)
@@
$\rho({\bf x},X_{loc}):=\inf\Big{\{}\sigma({\bf x},X_{loc},{\bf w}):\,{\bf w}\in{\mathbb{R}}^{n_{loc}},\;Dp({\bf x})=\sum_{{\bf x}_{j}\in X_{loc}}w_{j}p({\bf x}_{j})\;\text{ for all }p\in\Pi^{d}_{q}\Big{\}}.$
Despite the additional factor $\sqrt{n_{loc}}$ in (24) in comparison to the $\ell_{1}$-minimal formulas, both types of formulas behave similarly in numerical experiments, see further details in Davydov and Schaback [2018].
@@
In the case of Laplace operator $\Delta f=\frac{\partial^{2}}{\partial x_{1}^{2}}+\cdots+\frac{\partial^{2}}{\partial x_{d}^{2}}$ it is often attractive to enforce the positivity of the weight vector ${\bf w}$ in (21), by requiring that ${\bf x}\in X_{loc}$ and $w_{j}\geq 0$ for all ${\bf x}_{j}\in X_{loc}\setminus\{{\bf x}\}$. (Then $w_{j}<0$ for ${\bf x}_{j}={\bf x}$ because of the exactness for constant.) This for example helps to ensure that system matrices of the resulting meshless finite difference method for the Poisson equation are M-matrices, with various benefits for the numerical solution, see Seibold [2008] and references therein. In [Davydov and Schaback, 2018, Section 4.3] an error bound is obtained for positive formulas for elliptic differential operators of second order.
We generate inequality constrained $\ell_{2}$-minimal formulas in the form (5) and (12) for the operators $D_{{\boldsymbol{\eta}}_{i}}({\bf x}_{i})$ and $\Delta_{\bf x}({\bf x}_{i})$ that minimize (23) with parameters $r,\gamma$ chosen to take into account the error bound (22) as follows. The algorithm depends on the prescribed bounds $n_{\min},n_{\max}$ for the size of the set of influence and the artificial viscosity parameter $\mu_{i}\geq 0$. We start by choosing an initial set of influence $X_{i}=X_{i}^{init}$ consisting of $|X_{i}^{init}|=n_{\min}$ nearest neighbors of ${\bf x}_{i}$ in $X$, including ${\bf x}_{i}$ itself. (We use $n_{\min}=10$ in the numerical experiments in the bivariate case $d=2$.)
@@
and $X_{i}$ is initialized as $X_{i}^{init}$ if $\mu_{i}=0$ or $X_{i}^{visc}$ otherwise, and repeatedly expanded to the set of $\lceil 1.2\,|X_{i}|\rceil$ nearest neighbors of ${\bf x}_{i}$ as long as (26) is infeasible. Note that $v_{ii}$ is negative as soon as $\mu_{i}>0$ because of (17) and since $v_{ij}$, ${\bf x}_{j}\in X_{i}$, cannot all be zero thanks to (13). Again, we terminate the process if the new size is greater than $n_{\max}$, meaning the algorithm exits with failure. In this case we repeat the same procedure as above after removing the inequality constraints in (26) and resetting $X_{i}$ to $X_{i}^{init}$ if $\mu_{i}=0$ or $X_{i}^{visc}$ otherwise. With $n_{\max}=100$ in the bivariate case this happened very rarely and only for certain interior nodes situated very close to the inflow boundary nodes in Example 1 below, see Section 6.1. We set $v_{ij}=0$ for all $j$ such that ${\bf x}_{j}\in X_{i}\setminus X_{i}^{visc}$. Note that the largest size of $X_{i}$ observed in the numerical experiments of Section 6 was 27 for both the viscosity term ($X_{i}^{visc}$) and the divergence term (final $X_{i}$) when using the nodes of the Cartesian grid or Halton points, and reaching at most 40 for the viscosity term and 48 for the divergence term in the case of random nodes of Example 2 in Section 6.2.